# BEAM: Bridging Physically-based Rendering and Gaussian Modeling for Relightable Volumetric Video

### v1

@@ 3.2 PBR Materials Decomposition
To efficiently assign a roughness value to each Gaussian, we utilize a generative model that offers significantly faster performance compared to inverse rendering methods. Specifically, we feed the canonical mesh and multi-view images into the Material Diffusion module of CLAY [79] to generate a roughness texture for the mesh. Each valid pixel of the texture is then mapped back to the world coordinate system, allowing us to assign each Gaussian the roughness value of its nearest pixel.
Figure 4: We present the results of different rendering techniques. Real-time rendering focuses on efficiency, while offline rendering delivers more realistic shadows and occlusion effects.
Figure 5: Gallery of our results. We present some real-time rendering results under HDRI settings, which deliver high-fidelity rendering of human performances across challenging motions and complex clothing textures.
@@
For AO and base color, we first estimate these two attributes in the input views for each frame to bake 2D material maps, and then optimize into the corresponding dense skin Gausssians.
@@
$\mathcal{A}(x)=\frac{1}{\pi}\int_{\Omega}V^{\text{env}}\left(x,\boldsymbol{\omega}_{i}\right)(\mathbf{n}\cdot\boldsymbol{\omega}_{i})\mathrm{d}\boldsymbol{\omega}_{i},$ (3)
@@
For the base color, we use a simplified Disney BRDF model [2] composed of a Lambertian diffuse term and a Cook-Torrance specular term [8], with the outgoing radiance $L_{o}$ being a linear combination of these two components. Since for dielectric materials the diffuse part is proportional to the base color while the specular part is independent of it, the rendering equation [33] can be written as:
$L_{o}\left(x,\boldsymbol{\omega}_{o}\right)=\rho(x)L_{o}^{D}\left(x,\boldsymbol{\omega}_{o}\right)+L_{o}^{S}\left(x,\boldsymbol{\omega}_{o}\right),$ (4)
@@
where $L_{o}\left(x,\boldsymbol{\omega}_{o}\right)$ is the outgoing radiance at $x$ in direction $\boldsymbol{\omega}_{o}$, computed by mapping image pixel colors to linear space. $\rho$ is the base color, $L_{o}^{D}$ is the diffuse part residue and $L_{o}^{S}$ is the specular part. To simplify the rendering equation, our computation of the base color disregards the indirect illumination effects caused by surface reflections on the human body. Consequently, $L_{o}^{D}$ and $L_{o}^{S}$ are expressed as follows:
$L_{o}^{D}\left(x,\boldsymbol{\omega}_{o}\right)$ $=\frac{1}{\pi}\int_{\Omega}\left(1-F\right)\mathcal{L}(x,\boldsymbol{\omega}_{i})\mathrm{d}\boldsymbol{\omega}_{i},$ (5)
$L_{o}^{S}\left(x,\boldsymbol{\omega}_{o}\right)$ $=\int_{\Omega}f_{r_{s}}\left(x,\boldsymbol{\omega}_{i},\boldsymbol{\omega}_{o}\right)\mathcal{L}(x,\boldsymbol{\omega}_{i})\mathrm{d}\boldsymbol{\omega}_{i},$ (6)
@@
where $\mathcal{L}(x,\boldsymbol{\omega}_{i})=V^{\text{env}}\left(x,\boldsymbol{\omega}_{i}\right)L_{i}^{\text{env}}\left(\boldsymbol{\omega}_{i}\right)(\mathbf{n}\cdot\boldsymbol{\omega}_{i})$, $F$ is the approximated Fresnel term, $f_{r_{s}}$ is the specular term in the BRDF, $L_{i}^{\text{env}}$ can be queried from the environment map. To capture the environment map of our multi-view, well-lit dome setup, we position a DSLR camera at the center and take bracketed exposure photographs from multiple directions. These photographs are then processed using PTGui [47] to generate a high dynamic range (HDR) panoramic image, enabling precise calculation of the incoming radiance $L_{i}^{\text{env}}$ from the environment.
Figure 6: Qualitative comparisons of our method against R-3DGS [18], GS-IR [40] and D-2DGS [80]. D-2DGS refers to the results of inverse rendering applied to mesh sequences obtained from D-2DGS to decouple PBR materials. Our method achieves the highest relighting quality. For more detailed comparison results, please refer to the appendix.
@@
We observe that both 2D AO and base color need the visibility term $V^{\text{env}}$. To accurately compute $V^{\text{env}}$, we adopt Gaussian ray tracer from 3DGRT [45], and compute intersections based on the maximum Gaussian response. For efficient computation, a proxy icosahedron mesh is employed to leverage hardware acceleration, with a two-level BVH constructed at both the mesh and instance levels. We accumulate the alpha values of the Gaussians at the intersection points without sorting them:
$\alpha=\sum_{i=1}^{N}O_{i}G_{i}(x_{i}),$ (7)
@@
where $O_{i}$ is the opacity of the $i_{th}$ Gaussian, $G_{i}(x_{i})$ is the response of the Gaussian kernel at $x_{i}$ which makes the response be the maximum value along the ray. We set $V^{\text{env}}\left(x,\boldsymbol{\omega}_{i}\right)=0$ only when $\alpha$ is greater than a threshold $T$ (we choose $T$ as 0.9999); otherwise, $V^{\text{env}}\left(x,\boldsymbol{\omega}_{i}\right)=1$. Ray tracing stops once the accumulated alpha value exceeds the threshold $T$. To avoid occlusion from Gaussians directly above the original depth, we offset the ray origin $o$ along the normal direction: $o=x+\epsilon\mathbf{n}$.
@@
Finally, we use the Monte Carlo method to solve the integrals in Eq. 356. We render the AO image with 50 samples per pixel and the base color image with 100 samples per pixel, then apply Intel Open Image Denoise [25] for denoising.
@@
To optimize the AO and base color material maps into the corresponding dense skin Gaussians, we employ the generated training view material maps as ground truth supervision, while keeping all other geometry-related attributes fixed. The optimization process commences by initializing the AO attributes with zeros and the base color attributes with RGB values. We then train 5000 iterations to obtain both AO and base color attributes separately. Besides, we incorporate the same regularization term to the base color to ensure temporal consistency.
3.3 Physically Based Rendering
@@
By leveraging our relightable 4D Gaussians, we can seamlessly integrate the 4D assets into traditional CG engines, supporting both real-time and offline rendering workflows. For real-time rendering, we utilize the deferred shading to deliver immersive and efficient visualization across diverse settings. For offline rendering, we employ ray tracing, which excels in handling shadows and occlusion relationships, ensuring high-quality results.
@@
We implement real-time rendering using deferred shading techniques [12], based on the High Definition Render Pipeline (HDRP) in Unity. Specifically, we rasterize our 4D Gaussian sequence with extra PBR attributes, including base color, AO, normals, roughness, and depth maps, and store them in the GBuffer. This GBuffer is then integrated into the original forward transparent stage of HDRP for rendering semi-transparent objects. We also leverage shadow mapping from the HDRP rendering pipeline to perform shadow calculations for Gaussians under different light types. Our approach enables real-time rendering at 100 FPS in 1080P for volumetric videos.
@@
For offline rendering, we sort the intersection points of the Gaussian during ray tracing, and use alpha blending to determine the exact intersection with the entire Gaussian object, simultaneously acquiring the normal and material attributes at that point. This allows full compatibility with the widely used path tracing pipeline. To demonstrate the relighting quality of real-time rendering and offline rendering techniques, we showcase results for both under an environment map, as depicted in Fig. 4. Notably, the offline rendering effectively handles the shadows caused by occlusions in the Gaussian representations.
4 Experiments
@@
To demonstrate our relighting capabilities, we capture human performances with rich textures and challenge body movements using 81 calibrated industrial cameras, enabling high-quality rendering under diverse lighting conditions. Our pipeline is implemented based on 3DGS [35] and trained on a single NVIDIA GeForce RTX 3090. Our method achieves a processing time of 12 minutes per frame for 4D Gaussians modeling and optimization and 4 minutes for materials estimation and baking. The generated relightable 4D Gaussian sequences are fully compatible with VR platforms and CG engines, enhancing immersive experiences during playback and editing.
4.1 Comparison
We compare our method against state-of-the-art static Gaussian relighting techniques, R-3DGS [18], GS-IR [40], both of which reconstruct geometry frame by frame. For dynamic comparison, we first obtain the mesh sequences using D-2DGS [80] and then apply the path replay backpropagation method [63] for inverse rendering to decouple attributes, such as base color and roughness for each frame individually. As shown in Fig. 6, the normals and AO decoupled by Relightable-3DGS are blurry, resulting in significant relighting artifacts. GS-IR struggles to reconstruct smooth normals and effectively separate the base color from other attributes. While D-2DGS with inverse rendering produces smoother normals and more accurate attribute decoupling, the relighting results are prone to losing high-frequency details. In contrast, our method produces smooth normals and accurately decouples the AO and base color, enabling high-fidelity relighting results.
For quantitative comparison, we conduct evaluations on synthetic data to generate ground truth images under predefined lighting conditions. We use Blender to simulate the similar capture perspectives of our dome system and render human meshes from RenderPeople[51] into corresponding viewpoints using the CYCLES engine. we employ PSNR, SSIM, and LPIPS as metrics. To ensure a fair and precise comparison, we compute the metrics for AO, base color, and relighting results across two synthetic sequences, each consisting of 150 frames. In addition, we compute these metrics within the bounding box of the human region. As shown in Tab. 1, our method surpasses the other techniques in all metrics evaluated.
@@ 4.1 Comparison
Table 1: Quantitative comparison with SOTA relighting methods on our blender-rendered synthetic dataset. Green and yellow cell colors indicate the best and the second-best results.
@@
GS-IR 18.95 0.850 0.345 13.35 0.702 0.252 23.05 0.858 0.177
D-2DGS 24.19 0.906 0.232 21.04 0.839 0.145 23.23 0.857 0.198
Ours 25.32 0.924 0.168 21.47 0.906 0.084 25.57 0.895 0.086
Figure 7: Ablation on parameter settings for materials decomposition. Our full model provides accurate AO and base color representations, resulting in high-quality relighting with minimal artifacts.
Figure 8: Ablation study on the number of geometry-aware Gaussians. With $\sim 140,000$ Gaussians, BEAM achieves high rendering quality and geometric reconstruction accuracy, while ensuring efficient training and rendering.
4.2 Evaluation
@@
We conduct a qualitative ablation study on materials decomposition to the impact of different variables on AO, base color, and the final relighting results. For the offset variable in the Gaussians surface estimation during ray tracing, we present the results without the offset and with an excessively large offset (0.1), in the first and second columns of Fig. 7. We found that without an offset, AO and base color contained many black artifacts, and the Gaussian surface estimation suffered significant degradation. On the other hand, an offset of 0.1 led to overestimation, causing incorrect lighting decoupling and producing overly bright artifacts. Additionally, regarding the SH order required to represent AO and base color, as illustrated in the third column of Fig. 7, using a zero-order SH Gaussian resulted in blurriness. In contrast, our full ray tracing and material optimization pipeline, with an offset set to 0.02 and using third-order SH degree with geometry-aware Gaussians, provides accurate calculations and representations of the Gaussian’s AO and base color attributes.
@@
We evaluate the impact of varying the number of geometry-aware Gaussians on both rendering quality and mesh quality across 100 frames in our virtual dataset, using PSNR and Chamfer Distance respectively. As shown in Fig. 8, using $\sim 140,000$ geometry-aware Gaussians strikes an optimal balance between rendering quality and geometric reconstruction accuracy, without introducing excessive redundancy. This point count ensures efficient training and is well-suited for applications such as VR and AR.
4.3 User Study
We conduct a user study to evaluate the temporal reconstruction quality and normal consistency of our 4D Gaussians. We assess our method on our captured dataset, comparing it against R-3DGS, GS-IR, as well as D-2DGS, and ask 30 users to select the best option. For GS-IR and R-3DGS, we train Gaussian models for 200 frames individually and concatenate the per-frame normal maps. For our method and D-2DGS, we rasterize a normal map sequence from a 200-frame dynamic Gaussian sequence. In terms of temporal reconstruction quality, 95.65% of users prefer our method, while 87% choose our approach for normal consistency. These preference results clearly indicate a significant advantage of our method over the competing approaches, demonstrating its superior performance in both aspects.
5 Conclusion

### v2

@@ 3.2. PBR Materials Decomposition
To efficiently assign a roughness value to each Gaussian, we utilize a generative model that offers significantly faster performance compared to inverse rendering methods. Specifically, we feed the canonical mesh obtained from geometric optimization and multi-view images into the Material Diffusion module of CLAY (Zhang et al., 2024b) to generate a roughness texture for the mesh. Each valid pixel of the texture is then mapped back to the world coordinate system, allowing us to assign each Gaussian the roughness value of its nearest pixel.
@@
For AO and base color, we first estimate these two attributes in the input views for each frame to bake 2D material maps, and then optimize into the corresponding dense skin Gaussians.
@@
(4) $\mathcal{A}(x)=\frac{1}{\pi}\int_{\Omega}V^{\text{env}}\left(x,\boldsymbol{\omega}_{i}\right)(\mathbf{n}\cdot\boldsymbol{\omega}_{i})\mathrm{d}\boldsymbol{\omega}_{i},$
@@
For the base color, we use a simplified Disney BRDF model (Burley and Studios, 2012) composed of a Lambertian diffuse term and a Cook-Torrance specular term (Cook and Torrance, 1982), with the outgoing radiance $L_{o}$ being a linear combination of these two components. Since for dielectric materials the diffuse part is proportional to the base color while the specular part is independent of it, the rendering equation (Kajiya, 1986) can be written as:
(5) $L_{o}\left(x,\boldsymbol{\omega}_{o}\right)=\rho(x)L_{o}^{D}\left(x,\boldsymbol{\omega}_{o}\right)+L_{o}^{S}\left(x,\boldsymbol{\omega}_{o}\right),$
@@
where $L_{o}\left(x,\boldsymbol{\omega}_{o}\right)$ is the outgoing radiance at $x$ in direction $\boldsymbol{\omega}_{o}$, computed by mapping image pixel colors to linear space. $\rho$ is the base color, $L_{o}^{D}$ is the diffuse part residue and $L_{o}^{S}$ is the specular part. To simplify the rendering equation, our computation of the base color disregards the indirect illumination effects caused by surface reflections on the human body, and assumes the environment illumination is distant. Consequently, $L_{o}^{D}$ and $L_{o}^{S}$ are expressed as follows:
(6) $L_{o}^{D}\left(x,\boldsymbol{\omega}_{o}\right)$ $=\frac{1}{\pi}\int_{\Omega}\left(1-F\right)\mathcal{L}(x,\boldsymbol{\omega}_{i})\mathrm{d}\boldsymbol{\omega}_{i},$
(7) $L_{o}^{S}\left(x,\boldsymbol{\omega}_{o}\right)$ $=\int_{\Omega}f_{r_{s}}\left(x,\boldsymbol{\omega}_{i},\boldsymbol{\omega}_{o}\right)\mathcal{L}(x,\boldsymbol{\omega}_{i})\mathrm{d}\boldsymbol{\omega}_{i},$
@@
where $\mathcal{L}(x,\boldsymbol{\omega}_{i})=V^{\text{env}}\left(x,\boldsymbol{\omega}_{i}\right)L_{i}^{\text{env}}\left(\boldsymbol{\omega}_{i}\right)(\mathbf{n}\cdot\boldsymbol{\omega}_{i})$, $F$ is the approximated Fresnel term, $f_{r_{s}}$ is the specular term in the BRDF, $L_{i}^{\text{env}}$ can be queried from the environment map. To capture the environment map of our multi-view, well-lit dome setup, we position a DSLR camera at the center and take bracketed exposure photographs from multiple directions. These photographs are then processed using PTGui (New House Internet Services B.V., 2025) to generate a high dynamic range (HDR) panoramic image, enabling precise calculation of the incoming radiance $L_{i}^{\text{env}}$ from the environment.
@@
We observe that both 2D AO and base color need the visibility term $V^{\text{env}}$. To accurately compute $V^{\text{env}}$, we adopt Gaussian ray tracer from 3DGRT (Moenne-Loccoz et al., 2024) and compute intersections based on the maximum Gaussian response. For efficient computation, a proxy icosahedron mesh is employed to leverage hardware acceleration, with a two-level BVH constructed at both the mesh and instance levels. The $V^{\text{env}}$ can be computed from the Gaussians’ transmittance along the ray without sorting them:
(8) $V^{\text{env}}\left(x,\boldsymbol{\omega}_{i}\right)=\prod_{i=1}^{N}(1-o_{i}G_{i}(x_{i})),$
@@
where $o_{i}$ is the opacity of the $i^{th}$ Gaussian, $G_{i}(x_{i})$ is the response of the Gaussian kernel at the intersection point $x_{i}$. Ray tracing is early stopped if the attenuating $V^{\text{env}}$ is less than a threshold $T$ (we choose $T$ as 0.0001). To avoid occlusion from Gaussians directly above the original depth, we offset the ray origin $o$ along the normal direction: $o=x+\epsilon\mathbf{n}$.
Figure 4. We demonstrate scene and lighting editing within Unity, and immersive viewing using VR headsets.
Figure 5. We present the results of different rendering techniques. Real-time rendering focuses on efficiency, while offline rendering delivers more realistic shadows and occlusion effects.
@@
Finally, we use the Monte Carlo method to solve the $\mathcal{A}$ in Eq. 4, $L_{o}^{D}$ in Eq. 6, and $L_{o}^{S}$ in Eq. 7, then the base color $\rho$ can be solved from Eq. 5. For our Gaussian sampler, we render the AO image with 50 samples per pixel (spp) and the base color image with 100 samples per pixel, then apply Intel Open Image Denoise (Intel Corporation, 2025) for denoising.
@@
To optimize the AO and base color material maps into the corresponding dense skin Gaussians, we employ the generated training view material maps as ground truth supervision while keeping all other geometry-related attributes fixed. The optimization process commences by initializing the AO attributes with zeros and the base color attributes with RGB values. We then train 5000 iterations to obtain both AO and base color attributes separately. Besides, an $L_{2}$ regularization term is applied to the base color to ensure temporal consistency.
During optimization, the view-dependent RGB color $\boldsymbol{c}$ is used for material decomposition, but it is no longer required for relighting. After optimization, we discard this attribute and reparameterize each Gaussian as
(9) $\mathcal{P}^{\prime}_{i}(t)=\{\boldsymbol{\mu}_{i}(t),\boldsymbol{q}_{i}(t),\boldsymbol{s}_{i}(t),o_{i}(t),\boldsymbol{\rho}_{i}(t),r_{i}(t),\mathcal{A}_{i}(t)\}.\vskip-4.0pt$
This compact representation reduces storage and rendering load in CG engines and VR platforms, as $\boldsymbol{c}$ requires high-order spherical harmonics (SH) encoding to capture anisotropic effects.
3.3. Physically Based Rendering
@@
By leveraging our relightable 4D Gaussians, we can seamlessly integrate the 4D assets into traditional CG engines, supporting both real-time and offline rendering workflows. For real-time rendering, we utilize deferred shading to deliver immersive and efficient visualization across diverse settings. For offline rendering, we employ ray tracing, which excels in handling shadows and occlusion relationships, ensuring high-quality results.
@@
We implement real-time rendering using deferred shading techniques (Deering et al., 1988), based on the High Definition Render Pipeline (HDRP) in Unity. Specifically, we rasterize our 4D Gaussian sequence with extra PBR attributes, including base color, AO, normals, roughness, and depth maps, and store them in the GBuffer. This GBuffer is then integrated into the original forward transparent stage of HDRP for rendering semi-transparent objects. We also leverage shadow mapping from the HDRP rendering pipeline to perform shadow calculations for Gaussians under different light types. Our approach enables real-time rendering at 100 FPS in 1080P for volumetric videos.
Figure 6. Qualitative comparisons of our method against R-3DGS (Gao et al., 2025), GS-IR (Liang et al., 2024) and MeshAvatar (Chen et al., 2024). Our method achieves the highest relighting quality.
@@
For offline rendering, we sort the intersection points of the Gaussian during ray tracing and use alpha blending to determine the exact intersection with the entire Gaussian object, simultaneously acquiring the normal and material attributes at that point. This allows full compatibility with the widely used path tracing pipeline. To demonstrate the relighting quality of real-time rendering and offline rendering techniques, we showcase results for both under an environment map, as depicted in Fig. 5. Notably, the offline rendering effectively handles the shadows caused by occlusions in the Gaussian representations.
4. Application
We perform relighting experiments using the recovered geometry, material, and illumination from our BEAM pipeline. As shown in Fig. 3, the results demonstrate that our method effectively handles diverse lighting conditions, producing photorealistic renderings with accurate material response.
In Fig. 4, we showcase real-time editing of relightable 4D Gaussian sequences within the Unity platform, including lighting adjustment and scene composition. Our pipeline integrates smoothly with conventional CG workflows, enabling efficient scene and lighting editing for artists and designers. With the support of a custom Unity plugin, our system provides instant rendering feedback, significantly enhancing the speed and flexibility of iterative design and optimization.
Leveraging Unity’s cross-platform support, we seamlessly deploy the edited 4D scenes to VR headsets for immersive rendering and interaction. Multiple human performances can be aligned and rendered together in immersive environments, allowing for compelling and realistic virtual experiences. Users can explore the scene freely, interact with lighting and assets in real time, and experience performances from novel viewpoints, making it a powerful tool for virtual production, digital exhibitions, and immersive storytelling.
5. Experiments
@@
To demonstrate our relighting capabilities, we capture 6 diverse human performances featuring detailed textures and challenging body motions, using an array of 81 Z-CAM cinema cameras at 3840×2160 resolution and 30 FPS, under an adequately illuminated dome environment. Our pipeline is implemented based on 3DGS (Kerbl et al., 2023) and trained on a single NVIDIA GeForce RTX 3090 GPU. Our method achieves a processing time of 12 minutes per frame for 4D Gaussians modeling and optimization, and 4 minutes for materials estimation and baking. The generated relightable 4D Gaussian sequences are fully compatible with VR platforms and CG engines, enhancing immersive experiences during playback and editing.
5.1. Comparison
We compare our method with several state-of-the-art approaches. The static Gaussian relighting techniques, Relightable-3DGS (Gao et al., 2025) and GS-IR (Liang et al., 2024), reconstruct geometry frame by frame. The dynamic relighting method MeshAvatar (Chen et al., 2024) relies on an SMPL skeleton as a geometric proxy for lighting calculation. To ensure fairness, our comparison with MeshAvatar focuses on the relighting of reconstructed performance, rather than novel poses. We also provide these methods with the same environment maps used by our method. As shown in Fig. 6, the normals and AO decoupled by Relightable-3DGS are blurry, resulting in significant relighting artifacts. GS-IR struggles to reconstruct smooth normals and effectively separate the base color from other attributes. And MeshAvatar is constrained by mesh topology and SMPL prior, leading to reconstruction artifacts such as mesh distortion and surface tearing. These issues hinder the effectiveness of subsequent PBR-based disentanglement and relighting, resulting in suboptimal quality for high-frequency details. In contrast, our method produces smooth normals and accurately decouples the AO and base color, enabling high-fidelity relighting results.
@@ 5.1. Comparison
Table 1. Quantitative comparison with SOTA relighting methods on our synthetic dataset. Green and yellow cell colors indicate the best and the second-best results.
@@
GS-IR 18.95 0.850 0.345 18.86 0.773 0.281 23.05 0.858 0.177
MeshAvatar 19.53 0.735 0.459 21.04 0.867 0.146 23.93 0.857 0.178
Ours 25.32 0.924 0.168 21.47 0.906 0.084 26.57 0.895 0.086
For quantitative comparison, evaluations are conducted on synthetic data to generate ground truth images under predefined lighting conditions. We use Blender to simulate the similar capture perspectives of our dome system and render human meshes from RenderPeople(RenderPeople, 2025) into corresponding viewpoints using the CYCLES engine. Reconstruction quality is assessed using three widely used quantitative metrics: PSNR, SSIM, and LPIPS. To ensure a fair and precise comparison, we compute the metrics for AO, base color, and relighting results across two synthetic sequences, each consisting of 150 frames. In addition, we compute these metrics within the bounding box of the human region. As shown in Tab. 1, our method surpasses the other techniques in all metrics evaluated.
5.2. Evaluation
@@
We conduct a qualitative ablation study on materials decomposition to examine the impact of different variables on AO, base color, and the final relighting results. For the offset variable in the Gaussian surface estimation during ray tracing, we present the results without the offset and with an excessively large offset (0.1), in the first and second columns of Fig. 7. We observe that without an offset, AO and base color contained many black artifacts, and the Gaussian surface estimation suffered significant degradation. On the other hand, an offset of 0.1 led to overestimation, causing incorrect lighting decoupling and producing overly bright artifacts. Instead, an offset of 0.02 is applied for human data in the general case.
In addition, we evaluate the sampling strategy used in the Gaussian sampler for disentangling base color and AO. Specifically, we compare the effects of sampling count and the denoising component. We present results under three configurations: (1) Sample A: low sample counts (50 spp for AO, 100 spp for base color) without denoising, (2) Sample B: high sample counts (1000 spp for AO, 2000 spp for base color) without denoising, and (3) our adopted strategy using low sample counts with denoising. As shown in Fig. 8, our sampling strategy achieves a favorable balance between rendering quality and computation time, whereas alternative strategies either suffer from noisy outputs or incur significantly higher computational costs.
Figure 7. Ablation on offset variable for materials decomposition. Our full model provides accurate AO and base color representations.
Figure 8. Ablation on sampling strategies for material decomposition. Rendering time for each image is annotated in the zoomed-in views.
@@
We evaluate the impact of varying the number of 4D Gaussians on both rendering quality and mesh quality across 100 frames in our virtual dataset, using PSNR and Chamfer Distance, respectively. As shown in Fig. 9, we visualize how key metrics vary with the number of Gaussians, along with the mesh extracted from the Gaussian sequence at 140,000 points. We observe that using approximately 140,000 geometry-aware Gaussians achieves an optimal balance between rendering quality and geometric fidelity, while avoiding unnecessary redundancy. This point count ensures efficient training and is well-suited for applications such as VR and AR.
The Number of Cameras.
To assess our method’s robustness to sparse input views, we conduct an ablation study on the number of input camera views. While our proposed method was trained with 50 camera views on synthetic data, we also evaluate it on uniformly selected subsets of 20, 30, and 40 views. As shown in Tab. 2, our method demonstrates stable performance even with fewer cameras, showing its potential for generalization to sparse-view tasks.
Table 2. Quantitative ablation on the number of input camera views.
Methods Render Relighting
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
20 cams 34.56 0.935 0.0758 24.47 0.850 0.1056
30 cams 38.40 0.964 0.0598 25.07 0.854 0.0957
40 cams 39.73 0.975 0.0494 25.36 0.858 0.0924
Ours 42.73 0.980 0.0444 26.21 0.861 0.0892
5.3. User Study
We conduct a user study to evaluate the temporal reconstruction quality and normal consistency of our 4D Gaussians. For each method, we prepare visualizations consisting of normal maps and relighting results rendered under a consistent skybox environment to highlight illumination response. Specifically, for GS-IR and R-3DGS, Gaussian models are trained per frame over 200 frames, and both normal maps and relighting results are rendered under a shared HDR skybox. For our method and MeshAvatar, normal maps and relighting are generated from a dynamic 200-frame Gaussian sequence using the same skybox. We present the results to 30 users and ask them to select the most visually realistic rendering under novel lighting. In terms of temporal reconstruction quality and relighting quality, 95.65% of users prefer our method, while 87% choose our approach for normal consistency. These preference results clearly indicate a significant advantage of our method over the competing approaches, demonstrating its superior performance in both aspects.
Figure 9. Ablation study on the number of geometry-aware Gaussians. With $\sim 140,000$ Gaussians, BEAM achieves high rendering quality and geometric reconstruction accuracy, while ensuring efficient training and rendering.
6. Conclusion